# Optical and transport properties of NbN thin films revisited

### v2

@@ Abstract
Highly disordered NbN thin films exhibit promising superconducting and optical properties. Despite extensive study, discrepancies in its basic electronic properties persist. Analysis of the optical conductivity of disordered ultra-thin NbN films, obtained from spectroscopic ellipsometry by standard Drude-Lorentz model, provides inconsistent parameters. We argue that this discrepancy arises from neglecting the presence of quantum corrections to conductivity in the IR range. To resolve this matter, we suggest a modification to the Drude-Lorentz model, incorporating quantum corrections. The parameters obtained from the modified model are consistent with transport and superconducting measurements. The revisited values describing conduction electrons, which differ significantly from commonly adopted ones, are the electron relaxation rate Γ1.8eV/$\Gamma\approx 1.8~{}\textrm{eV}/\hbar$, the Fermi velocity $v_{F}\approx 0.7\times 10^{6}~{}\textrm{ms}^{-1}$ and the electron density of states $N(E_{F})=2~{}$states of both spins/eV/$V_{\textrm{f.u.}}$.
@@ II Optical properties
The suppression of $\sigma_{r}(\omega)$ revealed by mid-IR SE measurements results in a peak with maxima at $0.5-1$ eV. This ’anti-Drude’ behavior, also known as an anomalous or displaced Drude peak, was also predicted by the numerical study in Ref. 33 and extrapolation based on the KK relations.31 Thus, we attribute the decrease in the real part of the optical conductivity in the visible spectral range to the Drude-like $1/\omega^{2}$ behaviour typical for metals with finite relaxation rate $\Gamma$. As a consequence, $\Gamma$ in the studied NbN films is 1.5-2 eV.
@@
$\sigma_{QCD}(\omega)=\frac{\sigma_{0}}{1+(\frac{\omega}{\Gamma})^{2}}\left(1-\mathcal{Q}^{2}\left(1-\sqrt{\frac{\Omega}{\Gamma}}\right)e^{-\frac{1}{2}(\nicefrac{{\omega}}{{\Gamma}})^{2}}\right),$ (3)
where quantum correction exponentially vanishes at a scale $\Gamma/2$. This scale was fixed after its release during data fitting consistently yielded a uniform result. This may not be the case for much more disordered samples.32 The value of $\Omega$ is given by Eq. (2) and $T=300~{}$K, as the spectra were measured at room temperature.
@@
Here, $\mathcal{H}[\sigma_{QCD}(\omega)]$ denotes the Hilbert transform of the quantum-corrected Drude model, i.e. the transform of the first term of $\sigma_{r}(\omega)$ in Eq. (4), manifesting the Kramers-Kronig relations. The transform can be performed numerically, or by approximative analytical formula, which we derived in Appendix E. The second term of $\sigma_{i}$ is the imaginary part corresponding to the inter-band transition. This model of $\sigma_{r}(\omega)$ does not consider the inter-band transitions of bound electrons at high energies ($>7~{}$eV). Their effect on the imaginary conductivity is included in the third term in Eq. (5) via the parameter $\epsilon_{\infty}=1.6$ estimated in Appendix F.
The proposed model (thin lines in the Figs. 1 a) and b)) produces an excellent fit to both the real and the imaginary parts of the conductivity. Furthermore, utilizing only the visible range data the extrapolation by the fit to the proposed model is closely matched by the mid-IR optical conductivity (Fig. 1 b)), strongly supporting the model. The static conductivities were perfectly recovered by the fit, as well. The parameters of the best fit for the original and repeated analysis are listed in Table 1 and Table 5, respectively. The Drude conductivity $\sigma_{0}$ and the parameters of the inter-band transition peak exhibit variance, making it difficult to clearly observe their dependence on thickness. However, despite the variability in the scattering rate $\Gamma$, a subtle increasing trend with decreasing thickness can be inferred. A more pronounced thickness dependence is expected for thinner samples ($<6~{}$nm), as demonstrated, for instance, in Ref. 40. As $\Gamma$ rises, $\mathcal{Q}$ increases, too. This is consistent with expression for quantumness $\mathcal{Q}\approx 1/k_{F}l$, where $l=v_{F}/\Gamma$ and $v_{F}=\hbar k_{F}/m_{e}$ is the Fermi velocity, which we estimated as $v_{F}=\sqrt{\hbar\Gamma/(\mathcal{Q}m_{e})}\approx 0.7-0.8\times 10^{6}~{}$ms-1.26, 29
@@
[…]
@@ III Transport and superconducting properties
$d$ $L$ $\rho$ $R_{\square}$ $RRR$ $T_{c}$ $2k_{B}T_{c}$
[…]
@@
$1/\mathcal{Q}$ $\frac{\hbar}{3\mathcal{Q}m_{e}}$ $-\frac{4k_{B}}{\pi e}\left(\frac{\partial B_{c2}}{\partial T}\right)^{-1}_{T_{c}}$ $\sqrt{\frac{\hbar}{\mathcal{Q}\Gamma m_{e}}}$ $\sqrt{\frac{\hbar\Gamma}{\mathcal{Q}m_{e}}}$ $\frac{\hbar v_{F}}{\pi\Delta}$ $0.85\sqrt{\xi_{0}l}$ $\Phi_{0}/2\pi\xi_{GL}^{2}$ $\frac{3\sigma_{0}\mathcal{Q}m_{e}}{\hbar e^{2}}$ $\frac{\hbar\Gamma}{2\mathcal{Q}}$ $\frac{\sigma_{0}\Gamma m_{e}}{e^{2}}$
@@
The electron mean free path was estimated as $l=\sqrt{\hbar/\mathcal{Q}\Gamma m_{e}}\approx 2-3~{}$Å is slightly above half of the lattice parameter $a=4.4~{}$Å, which indicates that the samples are close to the Ioffe-Regel limit $k_{F}l\to 1$. The Ginzburg-Landau coherence length $\xi_{GL}(0)$ estimated by both Eqs. (9) and (11) ranges from 333~{}3nm to 444~{}4nm, which agrees with a commonly measured value in thin NbN films.54, 55, 45
@@ IV Discussion
It is now convenient to compare the revisited parameters of NbN with DFT simulations. We estimated the density of carriers in NbN as $n=\sigma_{0}\Gamma m_{e}/e^{2}=9\times 10^{28}~{}m^{-3}=2~{}V^{-1}_{\textrm{f.u.}}$, where $V_{\textrm{f.u.}}$ is the formula unit volume. For $\delta$-NbN, $V_{\textrm{f.u.}}=a^{3}/4$, where $a=4.4~{}\textrm{\AA}$ is the lattice constant. This result is in excellent agreement with the value obtained by integrating the DFT DOS from the threshold of the peak responsible for the modeled inter-band transition ($\approx E_{F}-4~{}$eV) up to $E_{F}$.42, 56, 37, 57, 58, 59 This also contradicts the common assumption, that all 4 Nb d-orbital electrons are conducting.17 We should emphasize here, that the agreement of our estimation of the number of electrons in the conductive band with the DFT result was obtained despite the fact that the DFT DOS itself is two times smaller than the ”Drude” DOS, which we determined from the Einstein relation $N(E_{F})=\sigma_{0}/(e^{2}D_{opt})$. This can be explained by the fact that the determined bottom of the conductive band $E_{c}=E_{F}-2~{}$eV is much closer to $E_{F}$ compared to DFT result 444~{}4eV. The value of $E_{F}-E_{c}$ was estimated from free electron relation between the density of electrons and DOS, given as $n=2N(E_{F})(E_{F}-E_{c})/3$. Similar effect was observed in ARPES measurement where the conductive band was significantly flattened in comparison to the DFT calculation, naturally leading to higher DOS at the Fermi level.41 The 111~{}1eV scale of smearing of the ARPES bandstructure also suggests a high value of scattering rate $\Gamma$. Nevertheless, ARPES agrees with DFT on the value of the Fermi momentum being approximately half of the $\Gamma$L path with length $\sqrt{3}\pi/a=1.24$-1. This agrees with the estimated $k_{F}\approx 0.6$-1.
For the ”Drude” DOS $N(E_{F})$, we obtained 2 states of both spins/eV/$V_{\textrm{f.u.}}$. Utilizing $\sigma_{DC}$ instead of $\sigma_{0}$, would lead to a different DOS, $N^{\prime}=(1-\mathcal{Q}^{2})N(E_{F})$ which is thickness-dependent.12, 13 The authors of Ref. 13 successfully compared this thickness-dependent $N^{\prime}$, obtained from transport measurements, to the DFT-calculated value, although the two are inherently different in nature. Furthermore, it is important to emphasize that DFT calculations only partially account for electron-electron and electron-phonon interactions, both of which are significant in NbN.60 Therefore, comparing the DFT value of 0.81.20.81.20.8–1.20.8 – 1.2 $\textrm{eV}^{-1}\textrm{V}_{\textrm{f.u.}}^{-1}$ with $N^{\prime}=0.5-1~{}\textrm{eV}^{-1}\textrm{V}_{\textrm{f.u.}}^{-1}$, or even with $N(E_{F})\approx 2~{}\textrm{eV}^{-1}\textrm{V}_{\textrm{f.u.}}^{-1}$, is questionable. Nevertheless, $N(E_{F})$ and $N^{\prime}$ can be considered upper and lower bound, respectively, for the actual DOS.
@@ Appendix A Mid-infrared range optical properties and samples aging
[…]
@@ Appendix C Effective medium theory
In Ref. 8, the double ENZ was explained via Maxwell-Garnet (MG) effective medium theory, which describes the optical properties of composite materials consisting of polarizable inclusions in an insulating matrix. In the case of NbN, NbN nanoparticles are immersed in a matrix of insulating niobium oxides. Taking the dielectric function of the oxide from Ref. 62, the resulting MG formula provides $\epsilon(\omega)$ with two zeros. However, such effective medium with an insulating matrix is inevitably an insulator. Moreover, the presence of oxygen in the NbN films was associated with the degree of porosity and it manifests itself in a significant increase of the resistance and/or the residual-resistance ration (RRR).23 Considering the sheet resistance and RRR (see Table 2), our samples have negligible oxygen content. The presence of oxygen should affect the density of the films as well,23 which was estimated from XRR, to have a constant value of 7.87.87.8~{}7.8gcm33-3- 3, which is close the ideal cubic NbN value.
@@ Appendix D Drude-Smith model
This model fits our experimental data very well (see Fig. 5b). However, the resulting granule size is 555~{}5Å, which is at least one order of magnitude smaller than the grain size estimated by X-Ray diffraction (XRD) measurement. For reasonable parameters (the measured grain size $L\approx 10~{}$nm and relaxation rate $1/\tau=\Gamma\approx 2~{}\textrm{eV}/\hbar$), this model leads to a displaced Drude peak in THz frequency range (meV). Thus, the fit of this model to our data produces unreasonable parameters. We do not claim that this effect is not present as it still can play a role at much smaller energies, where our measurements are not sensitive.
@@
[…]
@@ Appendix E Approximative formula for imaginary part of the modelled conductivity
[…]
@@
$\alpha=$ $2-2^{7/4}\sin\Big{(}\frac{\pi}{8}\Big{)}+\frac{8}{\pi}\int_{0}^{\infty}\!\!\!\frac{x^{5/2}dx}{1+2x^{2}+2x^{4}}\approx 3.136,$ (29)
$\beta$ $=2^{5/4}\sin\Big{(}\frac{\pi}{8}\Big{)}-\frac{4}{\pi}\int_{0}^{\infty}\!\!\!\frac{x^{1/2}dx}{1+2x^{2}+2x^{4}}\approx 0.308.$
@@
[…]

### v3

@@ Abstract
[…]
@@ II Optical properties
The suppression of $\sigma_{r}(\omega)$ revealed by mid-IR SE measurements results in a peak with maxima at 0.510.5-1 eV. This ’anti-Drude’ behavior, also known as an anomalous or displaced Drude peak, was also predicted by the numerical study in Ref. 33 and extrapolation based on the KK relations.31 Thus, we attribute the decrease in the real part of the optical conductivity in the visible spectral range to the Drude-like $1/\omega^{2}$ behaviour typical for metals with finite relaxation rate $\Gamma$. As a consequence, $\Gamma$ in the studied NbN films is 1.5-2 eV.
@@
$\sigma_{QCD}(\omega)=\frac{\sigma_{0}}{1+(\frac{\omega}{\Gamma})^{2}}\left(1-\mathcal{Q}^{2}\left(1-\sqrt{\frac{\Omega}{\Gamma}}\right)e^{-2(\nicefrac{{\omega}}{{\Gamma}})^{2}}\right),$ (3)
where quantum correction exponentially vanishes at a scale $\Gamma/2$. This scale was fixed after its release during data fitting consistently yielded a uniform result. This may not be the case for much more disordered samples.32 The value of $\Omega$ is given by Eq. (2) and T=300T=300~K, as the spectra were measured at room temperature.
@@
Here, $\mathcal{H}[\sigma_{QCD}(\omega)]$ denotes the Hilbert transform of the quantum-corrected Drude model, i.e. the transform of the first term of $\sigma_{r}(\omega)$ in Eq. (4), manifesting the Kramers-Kronig relations. The transform can be performed numerically, or by approximative analytical formula, which we derived in Appendix E. The second term of $\sigma_{i}$ is the imaginary part corresponding to the inter-band transition. This model of $\sigma_{r}(\omega)$ does not consider the inter-band transitions of bound electrons at high energies (>7>7~eV). Their effect on the imaginary conductivity is included in the third term in Eq. (5) via the parameter $\epsilon_{\infty}=1.6$ estimated in Appendix F.
The proposed model (thin lines in the Figs. 1 a) and b)) produces an excellent fit to both the real and the imaginary parts of the conductivity. Furthermore, utilizing only the visible range data the extrapolation by the fit to the proposed model is closely matched by the mid-IR optical conductivity (Fig. 1 b)), strongly supporting the model. The static conductivities were perfectly recovered by the fit, as well. The parameters of the best fit for the original and repeated analysis are listed in Table 1 and Table 5, respectively. The Drude conductivity $\sigma_{0}$ and the parameters of the inter-band transition peak exhibit variance, making it difficult to clearly observe their dependence on thickness. However, despite the variability in the scattering rate $\Gamma$, a subtle increasing trend with decreasing thickness can be inferred. A more pronounced thickness dependence is expected for thinner samples (<6<6~nm), as demonstrated, for instance, in Ref. 40. As $\Gamma$ rises, $\mathcal{Q}$ increases, too. This is consistent with expression for quantumness $\mathcal{Q}\approx 1/k_{F}l$, where $l=v_{F}/\Gamma$ and $v_{F}=\hbar k_{F}/m_{e}$ is the Fermi velocity, which we estimated as $v_{F}=\sqrt{\hbar\Gamma/(\mathcal{Q}m_{e})}\approx 0.7-0.8\times 10^{6}$~ms-1.26, 29
@@
$(\textrm{nm})$ $(S\mu\textrm{m}^{-1})$ $(\textrm{eV})$ $(1)$ $(S\mu\textrm{m}^{-1})$ $(\textrm{eV})$ $(\textrm{eV})$
@@ III Transport and superconducting properties
$d$ $L$ $\rho$ $_{\square} RR$RRRR $T_{c}$ $2k_{B}T_{c}$
$(\textrm{nm})$ $(\textrm{nm})$ $(\textrm{gcm}^{-3})$ $(\Omega)$ $(1)$ $(K)$ $(\textrm{meV})$
@@
$1/\mathcal{Q}$ $\frac{\hbar}{3\mathcal{Q}m_{e}}$ $\frac{4k_{B}}{\pi e}\left(\frac{\partial B_{c2}}{\partial T}\right)^{-1}_{T_{c}}$ $\sqrt{\frac{\hbar}{\mathcal{Q}\Gamma m_{e}}}$ $\sqrt{\frac{\hbar\Gamma}{\mathcal{Q}m_{e}}}$ $\frac{\hbar v_{F}}{\pi\Delta}$ $0.85\sqrt{\xi_{0}l}$ $\Phi_{0}/2\pi\xi_{GL}^{2}$ $\frac{3\sigma_{0}\mathcal{Q}m_{e}}{\hbar e^{2}}$ $\frac{\hbar\Gamma}{2\mathcal{Q}}$ $\frac{\sigma_{0}\Gamma m_{e}}{e^{2}}$
@@
The electron mean free path was estimated as $l=\sqrt{\hbar/\mathcal{Q}\Gamma m_{e}}\approx 2-3$~Å is slightly above half of the lattice parameter a=4.4a=4.4~Å, which indicates that the samples are close to the Ioffe-Regel limit $k_{F}l\to 1$. The Ginzburg-Landau coherence length $\xi_{GL}(0)$ estimated by both Eqs. (9) and (11) ranges from 33~nm to 44~nm, which agrees with a commonly measured value in thin NbN films.54, 55, 45
@@ IV Discussion
[…]
@@ Appendix A Mid-infrared range optical properties and samples aging
$(\textrm{nm})$ $(S\mu\textrm{m}^{-1})$ $(\textrm{eV})$ $(1)$ $(S\mu\textrm{m}^{-1})$ $(\textrm{eV})$ $(\textrm{eV})$
@@ Appendix C Effective medium theory
In Ref. 8, the double ENZ was explained via Maxwell-Garnet (MG) effective medium theory, which describes the optical properties of composite materials consisting of polarizable inclusions in an insulating matrix. In the case of NbN, NbN nanoparticles are immersed in a matrix of insulating niobium oxides. Taking the dielectric function of the oxide from Ref. 62, the resulting MG formula provides $\epsilon(\omega)$ with two zeros. However, such effective medium with an insulating matrix is inevitably an insulator. Moreover, the presence of oxygen in the NbN films was associated with the degree of porosity and it manifests itself in a significant increase of the resistance and/or the residual-resistance ration (RRR).23 Considering the sheet resistance and RRR (see Table 2), our samples have negligible oxygen content. The presence of oxygen should affect the density of the films as well,23 which was estimated from XRR, to have a constant value of 7.87.8~gcm3-3, which is close the ideal cubic NbN value.
@@ Appendix D Drude-Smith model
This model fits our experimental data very well (see Fig. 5b). However, the resulting granule size is 55~Å, which is at least one order of magnitude smaller than the grain size estimated by X-Ray diffraction (XRD) measurement. For reasonable parameters (the measured grain size $L\approx 10$~nm and relaxation rate $1/\tau=\Gamma\approx 2~\textrm{eV}/\hbar$), this model leads to a displaced Drude peak in THz frequency range (meV). Thus, the fit of this model to our data produces unreasonable parameters. We do not claim that this effect is not present as it still can play a role at much smaller energies, where our measurements are not sensitive.
@@
$(\textrm{nm})$ $(S\mu\textrm{m}^{-1})$ $(\textrm{eV})$ $(1)$ $(\textrm{eV})$ $(S\mu\textrm{m}^{-1})$ $(\textrm{eV})$ $(\textrm{eV})$
@@ Appendix E Approximative formula for imaginary part of the modelled conductivity
$\mathcal{H}[\tilde{\sigma}_{\textrm{QCD}}(x)]=\frac{2x}{\pi}\mathrm{P}.\mathrm{V}.\int_{0}^{\infty}\frac{\tilde{\sigma}_{\textrm{QCD}}(s)}{x^{2}-s^{2}}dx.$ (24)
@@
[…]
The comparison of Eq. (25) (red dashed lines) with the numerical result (green dashed lines) for various values of quantumness $\mathcal{Q}=0,\ 0.5,\ 0.75,\ 1$ is in Fig. 6a. One can see a excellent match, except for the value $\mathcal{Q}=1$, where a slight disagreement can be seen. This can be treated by taking the next term in the expansion (22). We calculated the Hilbert transform of the function
@@
[…]
@@
Figure 6: a) Black lines are the modified Drude conductivity $\sigma_{r}(x)$ with quantum corrections of various strength $\mathcal{Q}=0,\ 0.5,\ 0.75,\ 1$. Green dashed lines are their corresponding Kramers-Kronig images obtained numerically. Red dashed lines are given by the formula (25) for the Hilbert transforms of $\tilde{\sigma}_{r}(x)$. b) Green dashed lines are the same as in a), red dashed lines are plots of the formula (27) for $\mathcal{H}[\tilde{\tilde{\sigma}}_{r}(x)]$.